OpenStudy (anonymous):

Good evening all. I have a test tomorrow and I have been having the hardest time with questions like this one. Please help. "Four employees at a crayon factory are working on four separate machines. Cynthia's machine produces 100 less than twice the number of crayons as Zane's machine every hour. Louisa's machine produces 50 crayons more than Zane's machine every hour. Evan's machine produces 20 crayons less than Cynthia's machine every hour. How many crayons must Zane's machine produce every hour in order for Cynthia and Zane output to equal Louisa and Evan's output?"

OpenStudy (anonymous):

Start by writing general formulas for each sentence

OpenStudy (anonymous):

CM (cynthia's machine) = 100 - 2*ZM (Zane's Machine)

OpenStudy (anonymous):

LM = ZM + 50

OpenStudy (anonymous):

EM (Evan's Machine) = CM - 20

OpenStudy (anonymous):

so is it correct to also say thay EM = 2*ZM-100-20 ? or EM = (2+ZM-100)-20 or neither?

OpenStudy (anonymous):

Now read the question... I interpret it to mean, how many crayons must Zane's Machine be if CM + ZM = LM + EM

OpenStudy (anonymous):

1 second. I'm gonna get a pencil :P

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

I got a pencil... Give me a little bit. This question isn't worded very clearly

OpenStudy (anonymous):

ok thank you so much! I have a test on this and I am dying.

OpenStudy (anonymous):

First of all, this is a bad question.

OpenStudy (anonymous):

What you wrote is correct though. EM = 2*ZM - 100 - 20

OpenStudy (anonymous):

Since CM = 2*ZM - 100 (100 less than TWICE the output of Zane's machine)

OpenStudy (anonymous):

The reason I say this question is bad is because, I interpret the question to be asking, what is Zane's machine's output in order to make ZM + CM = LM + EM

OpenStudy (anonymous):

right, me too.

OpenStudy (anonymous):

But, after doing some algebra, it seems that is impossible since you will end up with 3 ZM on the left side of the equation and 3 ZM on the right side of the equation

OpenStudy (anonymous):

Are you sure you copied it correctly?

OpenStudy (anonymous):

I will recheck now.

OpenStudy (anonymous):

Hey... I'm in university... I have to get up for class tomorrow.

OpenStudy (anonymous):

sorry it took so long I wanted to upload a picture.

OpenStudy (anonymous):

Yea no worries

OpenStudy (anonymous):

This question is poorly worded then.

OpenStudy (anonymous):

If all of the variables (cynthia's output, louisa's output, and evan's output) are dependent upon lane's output, then we end up with an equation like follows:

OpenStudy (anonymous):

CM + ZM = EM + LM

OpenStudy (anonymous):

(2*ZM - 100) + ZM = (CM - 20) + (ZM + 50)

OpenStudy (anonymous):

(2*ZM - 100) + ZM = ((2*ZM - 100) - 20) + (ZM + 50)

OpenStudy (anonymous):

3ZM - 100 = 3ZM -70

OpenStudy (anonymous):

There's no solution

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Either the question has no answer or it is worded badly

OpenStudy (anonymous):

In my opinion...

OpenStudy (anonymous):

Well, I'm in university, I have class tomorrow (including a test) so I got to go

OpenStudy (anonymous):

Need some sleep. Talk to you later and good luck with the test.

OpenStudy (anonymous):

thank you so very much!

OpenStudy (anonymous):

np, hope it helped.

OpenStudy (anonymous):

it did, thanks!

OpenStudy (anonymous):

The trick to those problems it to write out each sentence as a mathematical statement.

OpenStudy (anonymous):

Take your time and build your facts.

OpenStudy (anonymous):

ok